# TE-NeXt: A LiDAR-Based 3D Sparse Convolutional Network for Traversability Estimation

### v1

@@ IV-A Datasets
The datasets contain around 25 to different semantic labels that a point can be assigned, such as bush, mud, asphalt, sidewalk, etc. The condition labels were reclassified based on its traversability, resulting in only two classes: ”traversable” and ”non-traversable”. The classes converted to ”traversable” include sidewalk, asphalt, low vegetation or grass, dirt, cement, and mud. The ”non-traversable” class comprises tree, person, car, truck, building, and others. This binary categorization was done by considering the ability of a robot to safely navigate across the different surface types.
@@ IV-B Training and evaluation process
Since the intention is to develop a multimodal model without giving preference to any of the environments described above, a prior selection has been carried out on the training to ensure equivalence. As Rellis-3D is not as extensive as SemanticKITTI, it was not used in its entirety. Specifically, 10,700 clouds of unstructured environments have been selected, belonging to 3 of the 4 sequences in Rellis-3D. The first 4 SemanticKITTI sequences have been selected, thus ensuring 11,100 clouds of structured environments. In addition, the validation and testing process have been carried out with sequences from both datasets, adding all the SemanticUSL data in order to assess the degree of generalisation of the neural network in a never-before-seen environment.
Table I specifies the hyperparameters and settings used to train the model from randomly initialized weights. The optimizer used AdamW with an initial Learning Rate (LR) of 5e-3 along to Cosine Annealing scheduler causes the LR to follow a cosine function over the same training epoch. It performs periodic ”warm restarts” where the learning rate is reset back to the initial value after a certain number of batches. This simulates the restarting of the learning process using the weights from the previous run as a good starting point. The periodic restarts help the model escape from local minima and saddle points. Since the model has been trained for a binary segmentation task, the F1 score is a suitable metric for imbalanced datasets and considers both: precision (the proportion of true positive predictions among all positive predictions) and recall (the proportion of true positive predictions among all actual positive instances). This ensures that the model is not optimizing for one metric at the expense of the other. Furthermore, if the F1 score on the validation set stops improving or starts to decrease, training can be stopped early to prevent overfitting and obtain a model with good generalization.
TABLE I: Environmental-independent training setting
@@ IV-C Ablation study
This section presents the roadmap followed during the development of the architecture proposed in this paper, being the starting point the MinkU-Net34C architecture. For this purpose, a first distinction will be made in three blocks. The first will provide a study of the effect of input features on network performance. The second one will present the design steps that affect the structure of the neural network. Subsequently, the methodology employed in the design of the residual block, which represents the primary innovation of this work, will be explained.
Out of many experiments, those successful were classified according to different design principle, as shown in Figure 4. Each bar represents the performance of the different models on the Rellis-3D dataset, in terms of two metrics considered relevant to the task of traversability estimation: F1 score and precision. They are considered relevant mainly because of the class imbalance and the importance of false negatives/positives. Precision measures the model’s ability to avoid matching non-traversable areas as traversable (avoid false positives), which is crucial for safety. Recall measures the ability of the model to detect all traversability areas (avoidance of false negatives), which is important for efficient navigation. The F1 metric combines precision and recall into a single score, allowing assesing the balance between the two of them.
@@ IV-C1 Input data
  • Features: In contrast to 2D convolutions, where the input and output tensors have a fixed number of feature dimensions, typically called channels (e.g. RGB channels for images), 3D sparse convolutions allow sparse tensors to have an arbitrary number of feature dimensions. Each non-zero element in the sparse tensor can have a feature vector of arbitrary length. Therefore, as a first step, a study of the possible most discriminating features for estimating traversability was carried out. Among the most representative are those shown in Figure 4 (experiments A, B, C, referred to in Table II). However, the vector of 1’s was chosen as the main one for future modifications because it is equally discriminant and at the time of inference it represents an improvement in efficiency.
@@ IV-C2 Global Design
  • Cardinality: This term was introduced by ResNext [47], and refers to the number of parallel branches within each block. In relation to this, numerous configurations have been evaluated. Firstly, the MinkU-Net34 network with (2,3,4,6,2,2,2,2,2) branches per block. In order to reduce the number of model parameters it was decided to test the MinkU-Net18 and MinkU-Net14. These smaller versions, especially MinkU-Net14, have significantly fewer branches in each convolutional block compared to MinkU-Net34. Fewer branches means fewer filters and therefore fewer trainable parameters in the model. In summary, by carefully exploring smaller architectures, it was possible to find very efficient MinkU-Net configurations that achieve a good compromise between model size and performance. It shows that it is often possible to substantially compress neural networks without sacrificing their ability to learn and generalise. This change does not represent an improvement in terms of performance but does represent an improvement in efficiency as it will be shown in Table VI.
  • Skip connections: Unlike the U-Net architecture [36], TeNext presents a variation in the connections that are made between the respective parts of the encoder and decoder. As it has already been seen in works such as [26, 27], it does not seem appropriate for the traversability estimation task to maintain the original connections proposed by U-Net. Thus, this change in the global design of the network allows to obtain a 5% improvement with respect to its predecessor, experiment E, and a 3% improvement with respect to the initial neural network F1 metric, MinkU-Net, experiment D and C.
  • Modifications in the residual block: Previously, the residual block was the original residual block of ResNetv1 [48], so the necessary modifications were made to update it to its second version, replacing the 2D convolution operations with linear convolutions. This modification is represented in the Figure 4 by the experiment G mentioned in the Table II and implies an increase of 2% in terms of F1.
Figure 4: Design process of the proposed architecture, TE-NeXt, starting from a base MinkU-Net34C. All experiments are detailed in Table II.
@@
TABLE II: Experimental modifications carried out during the design of the proposed architecture, for which performance is presented in Figure 4
Steps Experimental Modifications
A Features: $\vec{T}_{c}=(x_{i},y_{i},z_{i})$
B Features: $\vec{T}_{c}=(x_{i},y_{i},z_{i})+\vec{T}_{n}=(n_{x},n_{y},n_{z})$
C Features: One’s Vector
D Cardinality: (2,3,4,6,2,2,2,2) $\rightarrow$ (1,1,1,1,1,1,1,1)
E 4 skip connections $\rightarrow$ 3 skip connections
F 3 skip connections $\rightarrow$ 2 skip connections
@@ D Cardinality: (2,3,4,6,2,2,2,2) $\rightarrow$ (1,1,1,1,1,1,1,1)
H Fewer Activations
I Kernel Size: 5 $\rightarrow$ 7
J ReLu $\rightarrow$ GeLu
K Batch Norm $\rightarrow$ LayerNorm
L Drop Path
@@ IV-C3 Micro Design
  • Fewer activations: While activation functions after each convolutional layer is the most common approach, there are specific architectures and scenarios where selectively removing or moving them could make sense, like for computational efficiency or parameter reduction in mobile networks. However, this requires careful experimentation as haphazardly removing activations especially in deeper layers could hurt performance by limiting the non-linearity and representational power of the network.
@@
  • Kernel size: Traditional CNNs typically use small kernels, Transformer-based architectures and Transformer-inspired CNN designs often employ much larger kernels, from 7x7 and beyond. This allows them to capture longer-range dependencies and broader context, contributing to their strong performance on various vision tasks. Hence, as the Figure 4 shows, the experiment I referenced in the Table II leads to a 1% improvement with respect to F1 metric.
  • Normalization: Likewise to activation functions, it has become common practice to place operations that normalise the activations of each layer because of their multiple benefits in the learning process. However, BatchNorm [49] computes statistics (mean and variance) over the whole batch, so its effectiveness may decrease with small batch sizes. LayerNorm [50] avoids this problem by normalising the statistics per instance. This means it is not affected by batch size. In summary, LayerNorm’s robustness to small batch sizes, its applicability to recurrent and sequential models, and its consistency between training and inference, explain its superior performance as shown in experiment K in Figure 4.
@@
  • GeLu activation function. The ReLU has been the most popular activation function in deep learning due to its simplicity and efficiency. However, recent advanced Transformer models have started using the GELU instead. GELU is a smoother variant of ReLU that can enable better training of deep networks by reducing issues like vanishing gradients and allowing useful gradients for negative input values. The incorporation of this technique in the block resulted in a 2% increase in yield, which is reflected in experiment L.
@@
  • Regularization technique. Dropout and drop-path are regularization techniques used in training deep neural networks to prevent overfitting and to improve generalization performance. Dropout randomly drops out (sets to zero) units in a layer during training, while drop-path randomly drops out entire paths or connections between layers. This prevents units or paths from relying too much on each other, forcing them to individually be more robust.
@@ IV-D Results
In this section a performance comparison of the proposed method and models from the literature of the semantic segmentation problem will be presented. Aiming at providing a fair comparison, all models have been adapted to the traversability estimation problem, trained from scratch and validated under the same conditions and data. The original codes of the compared methods have been adapted for the study and had to be modified in order to establish the comparison.
The following comparison has been made with methods covering the same problem from different approaches: from a multi-layer perceptron approach, through the well-known 2D convolutional networks to the novel sparse convolutions that allow efficient handling of three-dimensional data. Since TE-NeXt has been developed based on the performance of modifications from the base architecture design, in unstructured environments, as discussed IV-C Section, the comparison has been divided into 3 different categories:
@@ IV-D1 Performance in structured environments
Table III presents the results of several semantic segmentation methods for the estimation of traversability in urban environments, using the validation sequence of the SemanticKITTI dataset. The proposed method TE-NeXt, achieves the best overall performance with a F1 score of 96.0%. This is supported by the nature of the metric itself as it combines precision and recall, taking into account both false positives and false negatives.
False positives in autonomous driving imply an inherent risk, not only to the vehicle itself but also to its environment and occupants. This critical aspect is reflected in metrics such as mIoU (mean Intersection over Union for traversable and non-traversable areas) and TNR (True Negative Rate), which assess the model’s ability to correctly identify non-traversable areas. The proposed TE-NeXt method achieves the same TNR as state-of-the-art techniques such as Cyl3D, PVKD and SalsaNeXt. These advances, although they may seem modest, have significant implications for safety by reducing the false positive rate in the segmentation of obstacles and inaccessible areas.
TABLE III: Traversability estimation performance of semantic segmentation methods in urban environments.
@@
Methods Voxel F1 score Accuracy mIOU TPR TNR
PointNet [51] - 93.50 93.70 88.1 96.50 91.10
PointNet++ [52] - 83.36 84.28 72.45 81.15 87.52
SqueezeSegv3 [53] - 96.5 97.6 94.9 96.6 98.1
SalsaNeXt [54] - 95.9 97.30 94.25 95.0 98.5
P-SVM [28] - 89.2 91.7 83.9 89.0 93.4
RN++ [55] - 94.0 95.4 90.7 93.2 96.8
Cyl3D [29] - 95.8 96.8 93.45 95.3 97.7
PVKD [56] - 95.9 96.9 93.65 95.6 97.7
MinkU-Net [38] 0.2 93.7 95.20 90.05 94.0 93.8
MinkU-Net[38]+ ConvNeXt [14] 0.2 94.6 95.2 90,85 97.6 94.4
TE-NeXt (ours) 0.2 96.0 96.3 92.90 94.7 97.7
IV-D2 Performance in natural environments
Table IV presents the results of various semantic segmentation methods for traversability estimation in natural environments using the Rellis-3D dataset. Unlike in urban environments, it is remarkable the good performance of the most classic method, PointNet, outperforming more modern methods in most of the metrics extracted.
The proposed architecture, TE-NeXt, exhibits the best overall performance, with an F1 score of 82.0%, an accuracy of 88.3% and an mIoU of 77.0% for the non-risk and potentially dangerous areas for the robot’s integrity. This outperforms other MinkU-Net-based approaches such as MinkU-Net+ConvNeXt, as well as methods that transform the input cloud given by a LiDAR to an image using projective geometry such as SqueezeSegV3 and SalsaNeXt. In addition to a high mIoU for traversable and non-traversable areas, it achieves a significant true positive rate (TPR) of 81.3%, indicating a good ability to correctly identify traversable areas. At the same time, it has a high true negative rate (TNR) of 91.8%, demonstrating its effectiveness in discarding non-traversable areas.
By contrast, compared to results in urban environments (Table III), the overall performance is lower for all methods in the Rellis-3D dataset. This poses that semantic segmentation for traversability estimation in natural environments is a more challenging task due to the higher variability and complexity of scenes. Since there are no references to the evaluation of some of the mentioned methods in unstructured environments, it has been decided not to include them in the following Tables.
TABLE IV: Traversability estimation performance of semantic segmentation methods in natural environments.
Natural envs. Sem. seg Rellis-3D
Methods Voxel F1 score Accuracy mIOU TPR TNR
PointNet [51] - 80.3 87.2 74.90 76.43 92.8
PointNet++ [52] - 58.9 50.2 32.1 45.6 67.0
SqueezeSegv3 [53] - 70.8 89.1 71.15 65.3 95.1
SalsaNeXt [54] - 73.5 90.0 73.30 68.9 95.5
MinkU-Net[38] 0.2 75.8 83.3 69.15 80.0 84.9
MinkU-Net[38] + ConvNeXt [14] 0.2 78.3 85.9 72.75 77.8 89.8
TE-NeXt (ours) 0.2 82.0 88.3 77.00 81.3 91.8
IV-D3 Performance in mixed environments
Regarding the performance of the proposed model, it is shown that it is comparable with the best methods in terms of F1 score, even though all the modifications made with a bias towards performance in natural environments. Methods such as SqueezeSegV3 and SalsaNext show very competitive performance, especially in terms of accuracy and TNR. They significantly outperform PointNet and PointNet++.
@@ IV-D3 Performance in mixed environments
Methods Voxel F1 score Accuracy mIOU TPR TNR
PointNet [51] - 90.2 90.1 82.0 94.9 85.7
PointNet++ [52] - 87.13 87.0 77.10 91.4 83.1
SqueezeSegv3 [53] - 91.5 95.6 89.4 97.9 98.25
SalsaNeXt [54] - 95.0 97.3 93.5 95.9 97.8
MinkU-Net[38] 0.2 94.3 94.4 89.35 98.1 91.0
MinkU-Net[38]+ ConvNeXt [14] 0.2 95.0 95.1 90.6 98.3 92.2
TE-NeXt (ours) 0.2 95.0 95.0 90.6 97.9 92.5

### v2

@@ IV-A Datasets
The datasets contain around 25 to different semantic labels that a point can be assigned, such as bush, mud, asphalt, sidewalk, etc. The condition labels were reclassified based on its traversability, resulting in only two classes: “traversable” and “non-traversable”. The classes converted to “traversable” include sidewalk, asphalt, low vegetation or grass, dirt, cement, and mud. The “non-traversable” class comprises tree, person, car, truck, building, and others. This binary categorization was done by considering the ability of a robot to safely navigate across the different surface types.
@@ IV-B Training and evaluation process
Given the considerable discrepancy in the quantity of labelled data available from urban and natural environments, being the latter the primary focus of this research, the objective was to achieve a more balanced distribution of data across these domains. In particular, 10,700 clouds of unstructured environments have been selected, belonging to four of the five sequences in Rellis-3D. Additionally, the initial four SemanticKITTI sequences have been selected, thus ensuring 11,100 clouds of structured environments. Furthermore, the validation and testing process have been conducted with sequences from both datasets, incorporating all the SemanticUSL data to assess the degree of generalisation of the neural network.
Table II specifies the hyperparameters and settings used to train the model from randomly initialized weights. The optimizer used AdamW with an initial Learning Rate (LR) of 5e-3 along to Cosine Annealing scheduler causes the LR to follow a cosine function over the same training epoch. It performs periodic “warm restarts” where the learning rate is reset back to the initial value after a certain number of batches. This simulates the restarting of the learning process using the weights from the previous run as a good starting point. The periodic restarts help the model escape from local minima and saddle points. Since the model has been trained for a binary segmentation task, the F1 score is a suitable metric for imbalanced datasets and considers both: precision (the proportion of true positive instances among all positive predictions) and recall (the proportion of true positive predictions among all actual positive instances). This guarantees that the model is not optimising for one metric at the expense of the other. Moreover, should the F1 score on the validation set cease to improve or begin to decline, training can be terminated prematurely to avert overfitting.
TABLE II: Hyperparameters selected in training process
@@
Figure 4: Design process of the proposed architecture, TE-NeXt, starting from a base MinkU-Net34C. All experiments are detailed in Table III.
@@ IV-C Ablation study
This section presents the roadmap followed during the development of the architecture proposed in this paper, with the objective of adapting the initial architectural approach used for indoor semantic segmentation, such as MinkU-Net34C, to outdoor environments, particularly natural environments. For this purpose, a first distinction will be made in three blocks.
The initial study examines the impact of input features and their spatial distribution on the performance of the neural network. The second one will present the design steps that affect the structure of the neural network. Subsequently, the methodology employed in the design of the residual block, which represents the primary innovation of this work, will be explained.
Figure 4 provides a graphical representation of the research process. During the development of the architecture, we have endeavoured to conduct experiments that are mutually related, thereby establishing a temporal and logical consistency in the nodes that comprise the graph. Consequently, the X-axis represents the design and research process, while the Y-axis represents the performance of the various models on the Rellis-3D dataset, in terms of one metric considered relevant to the task of traversability estimation: the F1 score. This metric is regarded as a valuable measure of model performance, particularly in situations where data sets are imbalanced. It offers a well-balanced assessment of a model’s precision and recall, making it an essential tool for evaluating model performance across a wide range of applications. Precision denotes the model’s capacity to avoid misclassifying non-traversable regions as traversable (i.e., preventing false positives), which is crucial for ensuring safety. Recall, on the other hand, represents the model’s ability to identify all traversable regions (i.e., avoiding false negatives), which is essential for efficient navigation. The F1 metric integrates precision and recall into a unified score, enabling an assessment of the equilibrium between the two.
@@ IV-C1 Input data
  • Data Discretisation Size: Discretization involves converting continuous spatial data into discrete coordinates, storing them efficiently. This process is essential for handling high-dimensional sparse tensors in 3D data processing tasks. Nevertheless, for tasks involving three-dimensional perception, the impact of quantization on performance is relatively insignificant when the quantization size is less than 5 cm. Thus, the experiments designated as I.1, I.2 and I.3 have been identified as a preliminary phase in a series of subsequent experiments. A finer discretization captures more information but can be computationally expensive, while coarser discretization simplifies the data at the risk of losing important features. The choice of discretization parameter should balance these factors based on the specific requirements of the task and available resources.
  • Features: In contrast to 2D convolutions, where the input and output tensors have a fixed number of feature dimensions, typically called channels (e.g. RGB channels for images), 3D sparse convolutions allow sparse tensors to have an arbitrary number of feature dimensions. Each non-zero element in the sparse tensor can have a feature vector of arbitrary length. Therefore, a study of the most discriminating features for estimating traversability was carried out. Among the most representative are those shown in Figure 4 (experiments I.4, I.5, and I.6, referred to in Table III). Nevertheless, the vector of 1’s (initialising the feature associated with each point to a value of 1) was selected as the primary candidate for future modifications, as it is equally discriminant and, at the time of inference, represents an improvement in efficiency.
@@ IV-C2 Global Design
  • Cardinality: This term was introduced by ResNext [48], and refers to the number of parallel branches within each block. Related to this, a number of different configurations have been assessed. The initial configuration employed was the MinkU-Net34 network, comprising (2, 3, 4, 6, 2, 2, 2, 2, 2) branches per block. In order to reduce the number of model parameters, it was decided to test the MinkU-Net18 and MinkU-Net14, experiments G.1 and G.2. The smaller versions, in particular MinkU-Net14, exhibit a notable reduction in the number of branches per convolutional block in comparison to MinkU-Net34. This leads to a decrease in the number of filters and, consequently, in the number of trainable parameters in the model. In conclusion, through a meticulous examination of smaller architectural configurations, highly efficient MinkU-Net configurations have been identified that strike a balance between model size and performance. This illustrates that it is frequently feasible to significantly reduce the complexity of neural networks without compromising their capacity to learn and generalise. While this modification does not lead to an significant enhancement in performance, it does contribute to greater efficiency, as will be demonstrated in Table VII.
  • Skip connections: Unlike the U-Net architecture [37], TeNext presents a variation in the connections that are made between the respective parts of the encoder and decoder. In addition to the four skip connections that were previously implemented in the experiments, two and three connections corresponding to experiments G.3 and G.4 have been evaluated. Notwithstanding the fact that both exhibited a reduction of 2% in terms of the F1 metric, the more robust performance options, experiment G.2 (later mentioned as the alternative experimental branch) and experiment G.4 (a configuration comprising just two connections), were subjected to further investigation in subsequent experiments. This has been influenced by concepts from related research [23, 24].
  • Modifications in the residual block: The aforementioned configuration settings are based on the original convolution block of ResNetv1 [49]. Consequently, the experiments between G.5 and G.8 seek to examine the efficacy of alternative convolution blocks that are commonly employed in CNNs for semantic segmentation:
    • Squeeze-and-Excitation (SE) Block [50]
    • ConvNeXt Block [10]
    • BottleNeck Residual Block [38]
    • Inception-ResNetV2 Block [11]
    Of these, the Inception-ResNetv2 block is the most effective. Its distinctive feature is the integration of Inception’s parallel filter operations with residual connections, which serves to enhance feature learning. Furthermore, it employs a range of convolutional filter sizes, enabling the capture of diverse spatial features. Nevertheless, the results obtained in this experiment do not surpass those obtained by previous experiments utilising the original ResNetv1 block.
@@
TABLE III: Experimental modifications carried out during the design of the proposed architecture, for which performance is presented in Figure 4
Steps ID Experimental Modifications
Input
features
I.1
I.2
I.3
I.4
I.5
I.6
Discretization parameter: 0.05 cm
Discretization parameter: 0.2 cm
Discretization parameter: 0.5 cm
Features: $\vec{T}_{c}=(x_{i},y_{i},z_{i})$
Features: $\vec{T}_{c}=(x_{i},y_{i},z_{i},n_{x},n_{y},n_{z})$
Features initialized at ones: $\vec{T}_{c}=(1)$
Global
design
G.1
G.2
G.3
G.4
G.5
G.6
G.7
G.8
Cardinality: (2,3,4,6,2,2,2,2) $\rightarrow$ (2,2,2,2,2,2,2,2)
Cardinality: (2,3,4,6,2,2,2,2) $\rightarrow$ (1,1,1,1,1,1,1,1)
4 skip connections $\rightarrow$ 3 skip connections
3 skip connections $\rightarrow$ 2 skip connections
@@
[…]
Micro
design
M.1
M.2
M.3
M.4
M.5
M.6
M.7
Batch Normalization $\rightarrow$ Layer Normalization
Fewer activations
Kernel size 5 $\rightarrow$ Kernel size 3
Kernel size 5 $\rightarrow$ Kernel size 9
Kernel size 5 $\rightarrow$ Kernel size 7
ReLu $\rightarrow$ GeLu
Drop Path Layer
@@ IV-C3 Micro Design
  • Normalization: Likewise activation functions, it has become common practice to place operations that normalise the activations of each layer because of their multiple benefits in the learning process. As previously stated, two distinct approaches are examined in this phase of the investigation. Firstly, within the alternative experimental branch, which has its origins in node G.2 and yielded the most favourable results using the original residual block of ResNetv1, the BatchNorm normalisation layer has been substituted with the Norm layer. However, this modification has caused a reduction in performance, with results declining from 78.3% to 75.6%. Secondly, the deployment of this layer within the configuration delineated by node G.8 has led to an enhancement of 0.6%. It may be reasonably deduced from the evidence presented that the normalisation layer in question performs more effectively after linear convolutions than after two-dimensional convolutions.
@@
  • Fewer activations: While activation functions after each convolutional layer is the most common approach, there are specific architectures and scenarios where selectively removing or moving them could make sense, like for computational efficiency or parameter reduction in mobile networks. However, this requires careful experimentation as haphazardly removing activations especially in deeper layers may impair performance by constraining the network’s nonlinearity and representational capacity. In the alternative experimental branch (represented as the union of G.2 and M.1), the elimination of the ReLu functions within the convolutional block of Resnetv1 is associated with a 3.8% reduction in performance. However, in the main experimentation branch, the previous results are surpassed for the first time, with an increase of 1.4% in comparison to the previous configuration and 0.5% compared to the optimal result observed in the experiments. The global F1 score is 78.8%.
@@
  • Kernel size: Conventional CNNs typically utilise relatively small kernels. In contrast, Transformer-based architectures and CNN designs inspired by the Transformer often employ significantly larger kernels, with sizes beginning at 7x7 and beyond. This allows them to capture longer-range dependencies and broader context, contributing to their strong performance on various vision tasks. To explore the solutions proposed in Transformer-based architectures, a study of the convolution size has been conducted. Experiments M.3, M.4 and M.5, depicted in Figure 4 and detailed in Table III, demonstrate that a larger convolution window size leads to superior outcomes compared to the conventional convolution window sizes typically employed in convolutional networks (2, 3 or 5). Consequently, this variation results in an increase of 0.9% in the F1 metric with respect to the previous optimal result.
@@
  • GeLu activation function. The ReLU has been the most popular activation function in deep learning due to its simplicity and efficiency. However, recent Transformer models have started using the GELU instead. GELU is a smoother variant of ReLU that can enable better training of deep networks by reducing issues like vanishing gradients and allowing useful gradients for negative input values. This activation function, experiment M.6, has resulted in an enhancement in performance relative to its predecessor, with an F1 metric of 80.5%.
@@
  • Regularization technique. Dropout and drop-path are regularization techniques used in training deep neural networks to prevent overfitting and to improve generalization performance. Dropout randomly drops out (sets to zero) units in a layer during training, while drop-path randomly drops out entire paths or connections between layers. This prevents units or routes from becoming excessively reliant on one another, thereby ensuring their individual robustness. This results in a 1.5% increase, which represents a final outcome of 82% of the experimentation.
@@ IV-D Results
This section offers a comprehensive performance evaluation of the proposed method, comparing it with established semantic segmentation and traversability estimation models from the literature. To ensure a fair comparison, we focus on accessible semantic segmentation models and methods originally designed for traversability estimation. The analysis covers a diverse range of approaches, including multi-layer perceptrons, traditional 2D convolutional networks, and cutting-edge sparse convolutions for efficient 3D data processing. As emphasised throughout this paper, TE-NeXt is an architecture optimised for unstructured environments. However, it can also be regarded as a trade-off solution in urban environments. In order to facilitate a comprehensive and structured comparison, this analysis will be divided into three distinct sections, each of which will focus on a specific type of environment.
@@ IV-D1 Performance in structured environments
Table IV presents the results of several semantic segmentation methods for the estimation of traversability in urban environments, using the validation sequence of the SemanticKITTI dataset.
In the context of traversability estimation, our proposed solution, while effective, is outperformed by methods like SqueezeSegv3 and SalsaNext when assessed in a 2D domain using image projections of 3D data (rows 3 and 5 of Table IV respectively). However, this 2D evaluation process introduces significant challenges and limitations. The projection from three-dimensional (3D) to two-dimensional (2D) space inevitably leads to information loss due to occlusions and depth ambiguities, where multiple points may map to a single pixel. When projecting segmentation results back to 3D space, issues such as one-to-many mapping and interpolation errors at object boundaries emerge, which result in a noticeable decrease in performance metrics [51, 52].
Thus, these methods have also been evaluated in a 3D context in order to compare their performance under the same conditions as the other methods (rows 4 and 6 of Table IV respectively). A decrease in their performance is observed, which confirms the limitations of 2D to 3D projections. To illustrate, SqueezeSegv3 demonstrates a decrease in its F1 score from 96.5% in 2D to 95.8% in 3D, whereas SalsaNext exhibits a reduction from 95.9% to 94.9%. Consequently, while these methods may show superior results in 2D evaluations, they may not be optimal for critical traversability tasks in 3D real-world environments. In contrast, TE-NeXt demonstrates a solid balance between accuracy, precision, and reliability, achieving an F1 score of 96.0% in structured environments. Thus, making it more suitable and robust for semantic segmentation in challenging, 3D real-world scenarios.
The comparison presented in Table IV has been performed by training the semantic segmentation models from scratch, with the exception of Cyl3D, PVKD and P-SVM, which were adapted to the traversability problem in [25] using the same approach. Furthermore, it has included additional traversability estimation methods, such as ProtoNet [53], MPTI [54] and PBG [20]. These metrics have been collected from the manuscripts of the aforementioned approaches, which have been subjected to the same testing conditions and code is not accessible.
TABLE IV: Traversability estimation performance of semantic segmentation methods in urban environments.
@@
Methods 2D 3D F1 score Accuracy mIOU TPR TNR
PointNet [55] $\checkmark$ 93.5 93.7 88.1 96.5 91.1
PointNet++ [56] $\checkmark$ 83.3 84.2 72.4 81.1 87.5
SqueezeSegv3 [57] $\checkmark$ 96.5 97.6 94.9 96.6 98.1
SqueezeSegv3 [57] $\checkmark$ 95.8 96.2 92.7 95.7 96.6
SalsaNext [52] $\checkmark$ 95.9 97.3 94.2 95.0 98.5
SalsaNext [52] $\checkmark$ 94.9 95.5 91.2 93.9 96.7
P-SVM [25] $\checkmark$ 89.2 91.7 83.9 89.0 93.4
RN++ [58] $\checkmark$ 94.0 95.4 90.7 93.2 96.8
Cyl3D [26] $\checkmark$ 95.8 96.8 93.4 95.3 97.7
PVKD [59] $\checkmark$ 95.9 96.9 93.6 95.6 97.7
MinkU-Net [39] $\checkmark$ 93.7 95.2 90.0 94.0 93.8
MinkU-Net[39]+ ConvNeXt [10] $\checkmark$ 94.6 95.2 90.8 97.6 94.4
ProtoNet [53] $\checkmark$ - - 80.4 - -
MPTI [54] $\checkmark$ - - 85.9 - -
PBG [20] - $\checkmark$ - - 84.1 - -
TE-NeXt (ours) $\checkmark$ 96.0 96.3 92.9 94.7 97.7
@@
Methods 2D 3D F1 score Accuracy mIOU TPR TNR
PointNet [55] $\checkmark$ 90.2 90.1 82.0 94.9 85.7
PointNet++ [56] $\checkmark$ 87.1 87.0 77.1 91.4 83.1
SqueezeSegv3 [57] $\checkmark$ 94.6 97.1 93.0 94.2 98.2
SqueezeSegv3 [57] $\checkmark$ 93.8 94.3 89.1 91.6 96.7
SalsaNext [52] $\checkmark$ 95.0 97.3 93.5 95.9 97.8
SalsaNext [52] $\checkmark$ 94.0 94.7 90.2 94.4 97.8
MinkU-Net[39] $\checkmark$ 94.3 94.4 89.3 98.1 91.0
MinkU-Net[39]+ ConvNeXt [10] $\checkmark$ 95.0 95.1 90.6 98.3 92.2
TE-NeXt (ours) $\checkmark$ 95.0 95.0 90.6 97.9 92.5
IV-D2 Performance in mixed environments
Regarding the performance of the proposed model, it is shown that it is comparable with the best methods in terms of F1 score, even though all the modifications made with a bias towards performance in natural environments. Methods such as SqueezeSegV3 and SalsaNext show very competitive performance, especially in terms of accuracy and TNR. As with structured environments, these methods demonstrate inferior performance in three-dimensional contexts.
IV-D3 Performance in natural environments
Table VI presents the results of various semantic segmentation methods for traversability estimation in natural environments using the Rellis-3D dataset. The comparison presented in Table VI is of particular relevance, given that natural environments represent the primary focus of the research. Unlike in urban environments, it is remarkable the good performance of the most classic method, PointNet, outperforming more modern methods in most of the metrics extracted.
The proposed architecture, TE-NeXt, exhibits the best overall performance, with an F1 score of 82.0%, an accuracy of 88.3% and an mIoU of 77.0% for the non-risk and potentially dangerous areas for the robot’s integrity. This outperforms other MinkU-Net-based approaches such as MinkU-Net+ConvNeXt, as well as techniques that transform the LiDAR input cloud into an image using projective geometry such as SqueezeSegV3 and SalsaNext.
In addition to a high mean intersection over union (mIoU) for traversability and non-traversability, the model achieves a true positive rate (TPR) of 81.3%, indicating a robust capacity to accurately identify traversability areas. Concurrently, it exhibits a high true negative rate (TNR) of 91.8%, which validates its efficacy in excluding non-traversability regions. In most of the metrics obtained, the TE-NeXt model demonstrates superior performance, although it is sometimes surpassed by the methods evaluated in two-dimensional space, which, as previously stated, exhibit significantly diminished performance when assessed through the same 3D criteria as the other methods.
In regard to methods dedicated to traversability estimation, such as nn Risk [60] and ScaTE [22], which is at the state of the art as evidenced in Table I, TE-NeXt is superior in the single available metric, as the code is not accessible.
Figure 5 presents the precision-recall curves for the evaluated traversability estimation methods, providing a comprehensive visualization of their performance across different classification thresholds. This graphical representation offers valuable insights into the trade-offs between precision and recall for each method, particularly in identifying non-traversable areas. As observed in the figure, the proposed TE-NeXt method (blue line) exhibits superior performance. It maintains high precision across a wide range of recall values, suggesting robust identification of non-traversable areas with minimal false positives. This characteristic is crucial for traversability estimation in autonomous navigation, where maintaining high precision (minimizing false positives) is often prioritized to ensure safety. The area under the precision-recall curve (AUC-PR), provides a quantitative measure for the overall performance comparison. Based on the visual inspection of the curves, TE-NeXt presents a high edge in this regard, suggesting its potential as a robust solution for traversability estimation in autonomous navigation systems.
TABLE VI: Traversability estimation performance of semantic segmentation methods in natural environments.
Natural envs. Sem. seg Rellis-3D
Methods 2D 3D F1 score Accuracy mIOU TPR TNR
PointNet [55] $\checkmark$ 80.3 87.2 74.9 76.4 92.8
PointNet++ [56] $\checkmark$ 58.9 50.2 32.1 45.6 67.0
SqueezeSegv3 [57] $\checkmark$ 70.8 89.1 71.1 65.3 95.1
SqueezeSegv3 [57] $\checkmark$ 70.2 85.5 68.1 67.8 94.1
SalsaNext [52] $\checkmark$ 73.5 90.0 73.3 68.9 95.5
SalsaNext [52] $\checkmark$ 70.9 85.0 70.5 65.4 95.0
MinkU-Net[39] $\checkmark$ 75.8 83.3 69.1 80.0 84.9
MinkU-Net[39] + ConvNeXt [10] $\checkmark$ 78.3 85.9 72.7 77.8 89.8
nn Risk [60] $\checkmark$ - - - 70.4 -
ScaTE [22] $\checkmark$ - - - 78.5 -
TE-NeXt (ours) $\checkmark$ 82.0 88.3 77.0 81.3 91.8
Figure 5: Performance comparison between methods trained from scratch by means of Precision-Recall curve on Rellis-3D dataset.